OpenStudy (anonymous):

Use the properties of exponents to simplify the following as much as possible 2x^6(−6x^5)

OpenStudy (anonymous):

=-12x^11 and if integrated then =-x^12 the simplest form..

OpenStudy (turingtest):

is this\[2x^6(-6x^5)\]?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

multiplication is commutative, meaning you can move the terms around, so this is the same as\[2x^6(-6x^5)=(2\cdot(-6))\cdot x^6\cdot x^5\]

OpenStudy (turingtest):

use the rule \[\large x^2\cdot x^b=x^{a+b}\] to simplify the variable terms